What is the point of intersection of | z-i |=4 and arg z=90^0?

Grade:11

1 Answers

arvind
8 years ago
arg(z)=90 implies the positive imaginary(Y) axis (equation....x=0)
| z-i |=4 implies a circle of radius 4 having centre at (0,i) …..equation{(x-0)^2 + (y-1)^2=16
solving both the equations
(y-1)^2=16
(y-1)=4
y=5...........note y is the modulus
therefore actual piont is(0,5i)